ICM 1998, Berlin, Aug. 18–27 Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: Sheshko, Michail, Marie Curie-Sklodowska University, Poland The direct method of solving of singular integral equations with Cauchy and Hilbert kernels